# On the kernels of the pro-$p$ outer Galois representations associated to once-punctured CM elliptic curves

### v2

@@ Abstract.
In this paper, we compare a certain field naturally arising from the kernel of the pro-$p$ outer Galois representation of a once-punctured elliptic curve with complex multiplication over an imaginary quadratic field $K$ with the maximal pro-$p$ Galois extension of the mod-$p$ ray class field $K(p)$ of $K$ unramified outside $p$. Among other things, we prove that the two fields coincide with each other for every split prime $p$ satisfying certain conditions, including an analogue of the so-called Deligne-Ihara conjecture. This result provides an analogue of Sharifi’s result on the kernel of the pro-$p$ outer Galois representation of the thrice-punctured projective line.
Keywords. once-punctured elliptic curve, complex multiplication, étale fundamental group, Galois representation.
Mathematics Subject Classification : 11G05, 11G15, 11R32, 12E30.
@@ 1. Introduction
In this paper, we study the kernel of the pro-$p$ outer Galois representations associated to once-punctured elliptic curves with complex multiplication.
@@
We recall the definition of the pro-$p$ outer Galois representations: Suppose that $X$ is a geometrically connected algebraic variety defined over a number field $F$. Throughout this paper, we work with a fixed algebraic closure $\bar{\mathbb{Q}}$ of $\mathbb{Q}$, and every number field is considered as a subfield of $\bar{\mathbb{Q}}$.
@@
We denote the étale fundamental group of $X$ with respect to a (possibly tangential) basepoint $\bar{x}$ by $\pi_{1}(X,\bar{x})$. In the following, we write $\bar{X}\coloneqq X\times_{F}\bar{\mathbb{Q}}$. There is an exact sequence determined by the structure morphism $X\to\mathrm{Spec}(F)$:
$1\to\pi_{1}(\bar{X},\bar{x})\to\pi_{1}(X,\bar{x})\to G_{F}\coloneqq\mathrm{Gal}(\bar{\mathbb{Q}}/F)\to 1$
@@
Hence the exact sequence, together with the conjugation action of $\pi_{1}(X,\bar{x})$ on $\pi_{1}(\bar{X},\bar{x})$, determines the outer Galois representation
$\rho_{X}\colon G_{F}\to\mathrm{Out}(\pi_{1}(\bar{X},\bar{x}))\coloneqq\mathrm{Aut}(\pi_{1}(\bar{X},\bar{x}))/\mathrm{Inn}(\pi_{1}(\bar{X},\bar{x}))$
@@
which does not depend on the choice of basepoints. For a prime $p$, since the maximal pro-$p$ quotient $\pi_{1}(\bar{X},\bar{x})^{(p)}$ of $\pi_{1}(\bar{X},\bar{x})$ is characteristic, the outer representation $\rho_{X}$ induces a homomorphism
$\rho_{X,p}\colon G_{F}\to\mathrm{Out}(\pi_{1}(\bar{X},\bar{x})^{(p)}),$
which we call the pro-$p$ outer Galois representation associated to $X$.
If $X$ is a hyperbolic curve, such outer representations are mainly studied in the context of anabelian geometry. Among other things, it is known that $\rho_{X}$ is injective if $X$ is a hyperbolic curve, cf. Matsumoto [Mat96] when $X$ is affine and Hoshi-Mochizuki [HM11] when $X$ is proper.
@@
As for the pro-$p$ outer Galois representation, it is far from being injective since the group $\mathrm{Out}(\pi_{1}(\bar{X},\bar{x})^{(p)})$ contains an open pro-$p$ subgroup. In particular, the fixed field of the kernel of $\rho_{X,p}$ is an almost pro-$p$ extension over $F$, and it seems to be interesting question to study arithmetic properties of this field.
@@
In the case of the thrice-punctured projective line $\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\}$, Anderson and Ihara [AI88, Theorem 2 (IV)] proved that the fixed field of the kernel of the pro-$p$ outer Galois representation is a nonabelian infinite pro-$p$ extension over the field $\mathbb{Q}(\mu_{p^{\infty}})$ unramified outside $p$, and is generated by certain $p$-units, which they call higher circular $p$-units, constructed in a certain geometric way. Moreover, they posed the following question [AI88, page 272, (a)]:
@@
Is ¯$\bar{\mathbb{Q}}^{\ker(\rho_{\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\},p})}$ the maximal pro-$p$ extension of $\mathbb{Q}(\mu_{p})$ unramified outside $p$?
@@
Let $\Omega^{\mathrm{cyc}}$ denote the maximal pro-$p$ extension of the $p$-th cyclotomic field $\mathbb{Q}(\mu_{p})$ unramified outside $p$. Regarding Anderson-Ihara’s question, Sharifi obtained the following affirmative answer for odd regular primes, assuming a certain conjecture which is nowadays referred to as the Deligne-Ihara conjecture:
Theorem (Sharifi [Sha02, Theorem 1.1]).
@@
Assume that $p$ is an odd regular prime and the Deligne-Ihara conjecture holds for $p$. Then the fixed field ¯$\bar{\mathbb{Q}}^{\ker(\rho_{\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\},p})}$ coincides with $\Omega^{\mathrm{cyc}}$.
@@
The Deligne-Ihara conjecture concerns with the structure of a graded Lie algebra over $\mathbb{Q}_{p}$ associated to a certain descending central filtration on $G_{\mathbb{Q}}$ induced by the pro-$p$ outer Galois representation of $\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\}$, and states that this graded Lie algebra is freely generated by Soulé elements $\{\sigma_{m}\}_{m\geq 3,\mathrm{odd}}$ in each odd degree $\geq 3$. We do not go into further details (see Sharifi [Sha02] and Ihara [Iha02] for detailed discussions). However, we will formulate an analogue of the Deligne-Ihara conjecture for once-punctured elliptic curves with complex multiplication (Conjecture 2.8).
@@
The Deligne-Ihara conjecture is now a theorem: The generation portion of the conjecture is proved by Hain-Matsumoto [HM03], and later Brown [Bro12] prove the conjecture in full generality. Combining Sharifi’s work with the settlement of the Deligne-Ihara conjecture, Anderson-Ihara’s question is affirmative if $p$ is odd and regular.
Remark 1.1.
@@
If $p$ is odd, the Galois group $\mathrm{Gal}(\Omega^{\mathrm{cyc}}/\mathbb{Q}(\mu_{p}))$ is generated by $(p+1)/2$ elements if and only if $p$ is regular by [NSW08, (10.7.13) Theorem]. This observation plays an essential role in the proof of [Sha02, Theorem 1.1]
If $p$ is odd and regular, the Galois group $\mathrm{Gal}(\Omega^{\mathrm{cyc}}/\mathbb{Q}(\mu_{p}))$ is even a free pro-$p$ group of rank $(p+1)/2$ by loc. cit. and we have an isomorphism
$\mathrm{Gal}(\Omega^{\mathrm{cyc}}/\mathbb{Q}(\mu_{p^{\infty}}))^{\mathrm{ab}}\cong\mathbb{Z}_{p}[[\mathrm{Gal}(\mathbb{Q}(\mu_{p^{\infty}})/\mathbb{Q}(\mu_{p}))]]^{\oplus\frac{p-1}{2}}$
between $\mathbb{Z}_{p}[[\mathrm{Gal}(\mathbb{Q}(\mu_{p^{\infty}})/\mathbb{Q}(\mu_{p}))]]$-modules by [NQD84, Corollaire 2.7], together with the fact that the left-hand side is generated by $(p-1)/2$ elements as an Iwasawa module.
Remark 1.2.
We remark that there exists a pair $(X,p)$ consisting of a hyperbolic curve $X$ over $\mathbb{Q}$ of positive genus and a prime $p$ such that the kernel of the pro-$p$ outer Galois representation $\rho_{X,p}$ coincides with that of the thrice-punctured projective line. For example, the Fermat curve $x^{p}+y^{p}+1=0$ of degree $p$ satisfies this condition [Hos15, Example 3.4 (i)]. In particular, for such a pair $(X,p)$, the fixed field of the kernel of $\rho_{X,p}$ coincides with $\Omega^{\mathrm{cyc}}$ if $p$ is odd and regular. In general, for every pair $(X,p)$ as above, it is known by Hoshi-Mochizuki [HM11, Theorem C (i)] that
$\ker(\rho_{X,p})\subset\ker(\rho_{\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\},p}).$
@@
The aim of this paper is to prove an analogue of Sharifi’s result in the case of once-punctured elliptic curves with complex multiplication.
@@
Let $K$ be an imaginary quadratic field of class number one, and $(E,O)$ an elliptic curve over $K$ with its origin $O\in E(K)$ having complex multiplication by the ring of integers $O_{K}$ of $K$. We assume that the complex multiplication by $O_{K}$ is defined over $K$, i.e. $O_{K}\hookrightarrow\mathrm{End}_{K}(E)$. In the following, for a prime $p$, we denote the mod-$p$ and mod-$p^{\infty}$ ray class fields of $K$ by $K(p)$ and $K(p^{\infty})$, respectively. The extension $K(p^{\infty})/K(p)$ is a $\mathbb{Z}_{p}^{2}$-extension.
@@
The following main result characterizes the kernel of the pro-$p$ outer Galois representation associated to the complement $X\coloneqq E-O$ of the origin of the elliptic curve, which we call the once-punctured elliptic curve.
Theorem A (Theorem 2.12).
Let $p\geq 5$ be a prime splitting into two distinct primes in $K$. Assume that
@@
    the class number of the mod-$p$ ray class field $K(p)$ is not divisible by $p$,
@@
    there are only two primes of the mod-$p^{\infty}$ ray class field $K(p^{\infty})$ lying above $p$, and
@@
Then the fixed field ¯$\bar{\mathbb{Q}}^{\ker(\rho_{X,p})}$ of the kernel of the pro-$p$ outer Galois representation of $X=E-O$ is equal to the compositum of the field $K(E[p])$ generated by the coordinates of $p$-torsion points of $E$ and the maximal pro-$p$ extension of $K(p)$ unramified outside $p$.
To the author’s knowledge, the pair $(X,p)$ as above gives a first example where the fixed field is determined independently of $(\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\},p)$, although it essentially depends on Conjecture 2.8.
In the following, we denote the maximal pro-$p$ extension of $K(p)$ unramified outside $p$ by $\Omega$. Note that $\Omega$ contains $\Omega^{\mathrm{cyc}}$, since $K(p)$ contains $\mathbb{Q}(\mu_{p})$.
Remark 1.3.
Let $p\geq 5$ be a prime which splits in $K$.
@@
    By [NSW08, (10.7.13) Theorem], we have
    $\dim_{\mathbb{F}_{p}}H^{1}(\mathrm{Gal}(\Omega/K(p)),\mathbb{F}_{p})=[K(p):K]+s_{p}+\dim_{\mathbb{F}_{p}}Cl_{p}(K(p))/p,$
    $\dim_{\mathbb{F}_{p}}H^{2}(\mathrm{Gal}(\Omega/K(p)),\mathbb{F}_{p})=-1+s_{p}+\dim_{\mathbb{F}_{p}}Cl_{p}(K(p))/p.$

### v3

@@ Abstract.
We prove that a certain field naturally arising from the kernel of the pro-$p$ outer Galois representation of a once-punctured elliptic curve with complex multiplication over an imaginary quadratic field $K$ coincides with the maximal pro-$p$ Galois extension of the mod-$p$ ray class field $K(p)$ of $K$ unramified outside $p$ under suitable assumptions. This result gives a geometric way to construct a large pro-$p$ extension of $K(p)$ with restricted ramification, and may be regarded as an analogue of Sharifi’s result in the case of the projective line minus three points.
Key words and phrases:
once-punctured elliptic curve, complex multiplication, étale fundamental group, Galois representation
1991 Mathematics Subject Classification:
11G05, 11G15, 11R32, 12E30
@@ 1. Introduction
In this paper, we study the pro-$p$ outer Galois representations associated to once-punctured elliptic curves with complex multiplication, especially their kernels. Let us briefly recall the definition of the pro-$p$ outer Galois representations: Suppose that $X$ is a geometrically connected algebraic variety defined over a number field $F$. Throughout the paper, we work with a fixed algebraic closure $\bar{\mathbb{Q}}$ of $\mathbb{Q}$, and every number field such as $F$ is considered to be a subfield of $\bar{\mathbb{Q}}$.
@@
We denote the étale fundamental group of $X$ with respect to a (possibly tangential) basepoint $\bar{x}$ by $\pi_{1}(X,\bar{x})$. In the following, we write $\bar{X}\coloneqq X\times_{\operatorname{Spec}(F)}\operatorname{Spec}(\bar{\mathbb{Q}})$. There is an exact sequence determined by the structure morphism $X\to\operatorname{Spec}(F)$, called the étale homotopy exact sequence:
$1\to\pi_{1}(\bar{X},\bar{x})\to\pi_{1}(X,\bar{x})\to G_{F}\coloneqq\operatorname{Gal}(\bar{\mathbb{Q}}/F)\to 1.$
@@
This exact sequence, together with the conjugation action of $\pi_{1}(X,\bar{x})$ on $\pi_{1}(\bar{X},\bar{x})$, determines the following outer Galois representation
$\rho_{X}\colon G_{F}\to\operatorname{Out}(\pi_{1}(\bar{X},\bar{x}))\coloneqq\operatorname{Aut}(\pi_{1}(\bar{X},\bar{x}))/\operatorname{Inn}(\pi_{1}(\bar{X},\bar{x})),$
@@
which does not depend on the choice of basepoints. For a prime $p$, the outer representation $\rho_{X}$ induces a homomorphism
$\rho_{X,p}\colon G_{F}\to\operatorname{Out}(\pi_{1}(\bar{X},\bar{x})^{(p)}),$
which we call the pro-$p$ outer Galois representation associated to $X$. Roughly speaking, since $\pi_{1}(\bar{X},\bar{x})^{(p)}$ is the fundamental group of the Galois category of geometrically $p$-coverings of $\bar{X}$, the outer representation $\rho_{X,p}$ expresses how $G_{F}$ acts on such a category or such coverings via pullbacks.
@@
If $X$ is a hyperbolic curve, these outer representations are studied in the context of anabelian geometry. Among other things, it is known that $\rho_{X}$ is injective if $X$ is a hyperbolic curve, cf. Matsumoto [Mat96] when $X$ is affine and Hoshi-Mochizuki [HM11] when $X$ is proper. Regarding the pro-$p$ outer Galois representation, it is far from being injective since the group $\operatorname{Out}(\pi_{1}(\bar{X},\bar{x})^{(p)})$ contains an open pro-$p$ subgroup. In particular, the fixed field of the kernel of $\rho_{X,p}$ is an almost pro-$p$ extension over $F$, and it seems to be interesting to study arithmetic properties of this fixed field.
@@
In the case of the thrice-punctured projective line $\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\}$, Anderson and Ihara [AI88, Theorem 2 (IV)] proved that the fixed field of the kernel of the pro-$p$ outer Galois representation is a nonabelian infinite pro-$p$ extension over the field $\mathbb{Q}(\mu_{p^{\infty}})$ unramified outside $p$, and is generated by certain $p$-units, which they call higher circular $p$-units, constructed from cusps of geometrically $p$-coverings of $\mathbb{P}^{1}_{\bar{\mathbb{Q}}}\setminus\{0,1,\infty\}$ of genus zero. Moreover, they posed the following question [AI88, page 272, (a)]:
@@
Is $\bar{\mathbb{Q}}^{\operatorname{Ker}(\rho_{\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\},p})}$ the maximal pro-$p$ extension of $\mathbb{Q}(\mu_{p})$ unramified outside $p$?
@@
Let $\Omega^{\mathrm{cyc}}$ denote the maximal pro-$p$ extension of the $p$-th cyclotomic field $\mathbb{Q}(\mu_{p})$ unramified outside $p$. Regarding Anderson-Ihara’s question, Sharifi [Sha02] has established the following affirmative result for odd regular primes, assuming a certain conjecture which is nowadays referred to as the Deligne-Ihara conjecture (it was called Deligne’s conjecture in [Sha02]):
Theorem 1.1 (Sharifi [Sha02, Theorem 1.1]).
@@
Let $p$ be an odd regular prime and suppose the Deligne-Ihara conjecture holds for $p$. Then the fixed field of the pro-$p$ outer Galois representation associated to the projective line minus three points coincides with $\Omega^{\mathrm{cyc}}$.
@@
The Deligne-Ihara conjecture concerns with the structure of a graded Lie algebra over $\mathbb{Q}_{p}$ associated to a certain descending central filtration on $G_{\mathbb{Q}}$ induced by the pro-$p$ outer Galois representation, and states that the graded Lie algebra is freely generated by Soulé elements $\{\sigma_{m}\}_{m\geq 3,\mathrm{odd}}$ in each odd degree $\geq 3$. We do not go into further details here (see Sharifi [Sha02] and Ihara [Iha02] for detailed discussions), but we will formulate an analogue of the Deligne-Ihara conjecture for once-punctured elliptic curves with complex multiplication in the next section (Conjecture 2.3).
@@
The Deligne-Ihara conjecture is now a theorem: The generation portion of the conjecture is proved by Hain-Matsumoto [HM03], and later Brown [Bro12] proved the conjecture in full generality. Anderson-Ihara’s question is hence affirmative whenever $p$ is odd and regular.
Remark 1.1.1.
@@
Let $p$ be an odd prime. The Galois group $\operatorname{Gal}(\Omega^{\mathrm{cyc}}/\mathbb{Q}(\mu_{p}))$ is generated by $(p+1)/2$ elements if and only if $p$ is regular by [NSW08, (10.7.13) Theorem]. This observation plays a central role in the proof of [Sha02, Theorem 1.1] Moreover, if $p$ is odd and regular, then $\operatorname{Gal}(\Omega^{\mathrm{cyc}}/\mathbb{Q}(\mu_{p}))$ is even a free pro-$p$ group of rank $(p+1)/2$ and the $p$-ramified Iwasawa module $\operatorname{Gal}(\Omega^{\mathrm{cyc}}/\mathbb{Q}(\mu_{p^{\infty}}))^{\mathrm{ab}}$ is generated by $\frac{p-1}{2}$ elements by loc. cit.. One can even show that the $p$-ramified Iwasawa module is free (of rank $\frac{p-1}{2}$) by using [NQD84, Corollaire 2.7].
@@
The aim of the present paper is to establish an analogue of Sharifi’s result in the case of once-punctured elliptic curves with complex multiplication defined over imaginary quadratic fields of class number one.
@@
In the following, let $K$ be an imaginary quadratic field of class number one, and $(E,O)$ an elliptic curve over $K$ with the origin $O\in E(K)$ having complex multiplication by the ring of integers $O_{K}$. We assume that the complex multiplication by $O_{K}$ is defined over $K$, i.e. $O_{K}\hookrightarrow\operatorname{End}_{K}(E)$. For a prime $p$ and a nonnegative integer $n$, we write $K(p^{n})$ for the ray class field of $K$ of conductor $p^{n}$, and set $K(p^{\infty})\coloneqq\cup_{n\geq 0}K(p^{n})$. Then $K(p^{\infty})/K(p)$ is a $\mathbb{Z}_{p}^{2}$-extension.
@@
Our main result characterizes the kernel of the pro-$p$ outer Galois representation associated to the complement $X\coloneqq E-O$ of the origin of the elliptic curve, which we call the once-punctured elliptic curve associated to $E$.
Theorem A (=Theorem 2.14).
Let $p\geq 5$ be a prime which splits in $K$. Assume that
@@
    the class number of $K(p)$ is not divisible by $p$,
@@
    there are only two primes of $K(p^{2})$ lying above $p$, and
@@
Then the fixed field of the kernel of the pro-$p$ outer Galois representation of X=EOX=E-O coincides with the compositum of $K(E[p])$ and the maximal pro-$p$ extension of $K(p)$ unramified outside $p$.
In the following, we write $\Omega$ for the maximal pro-$p$ extension of $K(p)$ unramified outside $p$. To the author’s knowledge, though Theorem A critically depends on our working Conjecture 2.3, the pair $(X,p)$ as above gives a first example where the fixed field is explicitly determined independently of the case of $(\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\},p)$. Note that one can construct many hyperbolic curves of higher genera whose kernels of pro-$p$ outer Galois representations are the same as that of the projective line minus three points by using ‘‘Oda’s prediction’’ [Tak12, Theorem 3.6] (or [HM11, Theorem C (i)]) and [Hos12, Lemma 28]. The Fermat curve of degree $p$ provides one such example [Hos15, Example 3.4 (2)].
Remark 1.1.2.
Let $p\geq 5$ be a prime which splits into distinct primes in $K$.
@@
    The first and second assumptions of Theorem A, together with [NSW08, (10.7.13) Theorem], imply that the pro-$p$ group $\operatorname{Gal}(\Omega/K(p))$ is generated by $[K(p):K]+2$ elements satisfying a single relation.
  2. (2)
    The first and second assumptions of Theorem A also allow us to determine the $\Lambda$-module structure of the $p$-ramified Iwasawa module $\operatorname{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}$ completely, where
    $\Lambda\coloneqq\mathbb{Z}_{p}[[\operatorname{Gal}(K(p^{\infty})/K(p))]]$
    is non-canonically isomorphic to the two-variable power series ring over $\mathbb{Z}_{p}$. More precisely, under these two assumptions, we have
    $\operatorname{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}\cong\Lambda^{[K(p):K]-1}\oplus\operatorname{Ann}_{\Lambda}(\mathbb{Z}_{p}(1)),$
    where $\operatorname{Ann}_{\Lambda}(\mathbb{Z}_{p}(1))$ denotes the annihilator ideal of $\mathbb{Z}_{p}(1)$, cf. Proposition 4.4.
  3. (3)
    We do not know whether there are infinitely many primes satisfying the first assumption of Theorem A, simply because the assumption implies that $p$ is regular. We also do not know whether the second assumption is satisfied by infinitely many primes. There are 39175 primes which are less than $10^{6}$ and split in $\mathbb{Q}(\sqrt{-1})$, and among them, 13705 primes satisfy the second assumption.
  4. (4)
    It is desirable to generalize Theorem A to more general imaginary quadratic fields. One might try to replace $X$ with a once-punctured CM elliptic curve over the Hilbert class field of the concerned imaginary quadratic field.
Example 1.2.
We have checked whether a small prime $p$ satisfies the first two assumptions of Theorem A by using SageMath [The24]. For example, when $p=5$, both assumptions are satisfied if and only if $K=\mathbb{Q}(\sqrt{-1})$. Similarly, when $p=7$, both conditions are satisfied only for $K=\mathbb{Q}(\sqrt{-3})$ or $\mathbb{Q}(\sqrt{-19})$. When $p=13$ or $17$, we have also checked that both assumptions are satisfied for $K=\mathbb{Q}(\sqrt{-1})$ under the Generalized Riemann Hypothesis (GRH).
Our strategy to prove Theorem A is to generalize Sharifi’s technique developed in [Sha02]. That is, we construct certain elements $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ that satisfy the assumption of Theorem A in such a group-theoretic way that they generate $\operatorname{Gal}(\Omega/K(p^{\infty}))$. Then Conjecture 2.3 allows us to conclude that the surjective homomorphism $\operatorname{Gal}(\Omega/K(p^{\infty}))\to\operatorname{Gal}(\Omega^{\ast}/K(p^{\infty}))$ is an isomorphism.
However, to construct such elements, we need to introduce a new ingredient. Namely, we define certain two-variable filtrations on the pro-$p$ geometric fundamental group $\Pi_{1,1}$ of $X$ and on the absolute Galois group $G_{K}$. We introduce them and establish their fundamental properties in Section 3.
As was mentioned in Remark 1.1.2 (1), in our situation, generators of the Galois group $\operatorname{Gal}(\Omega/K(p))$ satisfy a nontrivial relation. This is one of the significant differences from the previous result on the thrice-punctured projective line (cf. Remark 1.1.1), making it difficult to extend Sharifi’s approach to the case of once-punctured elliptic curves. We overcome this difficulty by choosing elements $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ carefully in Section 4. Once we obtain $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$, the rest of the proof follows in the same way as the previous result.
This paper proceeds as follows: In Section 2, we explain previous studies on the pro-$p$ outer Galois representations associated to once-punctured elliptic curves. Among other result, we review the construction of a certain power series obtained from the Galois action on the metabelian fundamental group and state Nakamura’s explicit formula for that power series. We also propose an analogue of the Deligne-Ihara conjecture and state the main result in a precise manner in that section. In Section 3, we introduce a two-variable version of the descending central series on various profinite groups and establish their fundamental properties. Section 4 is devoted to the proof of Theorem A, and in the appendix, we obtain a certain sufficient condition for the so-called pure locality for $K(\mu_{p})$, which is needed to establish the finiteness of certain Galois cohomology groups.
Acknowledgement. The present paper is based on a part of the author’s doctoral thesis submitted to Kyoto University. The author sincerely thanks his advisor Akio Tamagawa for helpful comments and warm encouragement, and Benjamin Collas for carefully reading drafts of the thesis. He also sincerely appreciates the anonymous referees for their constructive comments, which have greatly contributed to improving the content of the initial manuscript of the paper. The author was supported by JSPS KAKENHI Grant Number JP21J1191 when preparing his thesis, and by JP23KJ1882 throughout the writing of the the present paper.
Notational conventions
Indexes
Let $\boldsymbol{m}=(m_{1},m_{2})$ be a pair of integers and $w$ a positive integer. We write
  • $\boldsymbol{1}$ for the pair $(1,1)$,
  • $\lvert\boldsymbol{m}\rvert\coloneqq m_{1}+m_{2}$,
  • $\boldsymbol{m}\equiv 0\bmod w$ if both $m_{1}$ and $m_{2}$ are divisible by $w$.
Suppose $\boldsymbol{n}=(n_{1},n_{2})$ is another pair of integers. We write
  • $\boldsymbol{m}\geq\boldsymbol{n}$ if $m_{i}\geq n_{i}$ for $i=1,2$,
  • $\boldsymbol{m}>\boldsymbol{n}$ if $\boldsymbol{m}\geq\boldsymbol{n}$ and $\lvert\boldsymbol{m}\rvert>\lvert\boldsymbol{n}\rvert$.
Profinite Groups
Let $G$ be a profinite group, and $H,K$ two closed subgroups of $G$. We write
  • $[H,K]$ for the closure of the commutator subgroup of $H$ and $K$,
  • $\{G(m)\}_{m\geq 1}$ for the descending central series defined by
  • $G^{\mathrm{ab}}$ for the maximal abelian quotient $G/G(2)$,
  • $G^{\rm{mab}}$ for the maximal metabelian quotient $G/[G(2),G(2)]$ of $G$.
  • $G^{(p)}$ for the maximal pro-$p$ quotient of $G$ (for a prime $p$),
  • $\operatorname{Aut}(G)$ for the (continuous) automorphism group of $G$,
  • $\operatorname{Inn}(G)$ for the inner automorphism group of $G$,
  • $\operatorname{Out}(G)$ for the outer automorphism group $\operatorname{Aut}(G)/\operatorname{Inn}(G)$ of $G$.
Let $S$ be a subset of $G$. We write
  • $\langle S\rangle$ for the minimal closed subgroup of $G$ that contains $S$,
  • $\langle S\rangle_{{\rm normal}}$ for the minimal normal closed subgroup of $G$ that contains $S$.
Moreover, we say
  • $S$ generates $G$ if $\langle S\rangle$ coincides with $G$,
  • $S$ strongly generates $G$ if $S$ generates $G$ and converges to $1$ (cf. Ribes-Zalesskii [RZ10, 2.4]),
  • $S$ normally generates $G$ if $\langle S\rangle_{{\rm normal}}$ coincides with $G$.
Free pro-$p$ groups of countably infinite rank often appear in the paper. We refer to Ribes-Zalesskii [RZ10, Lemma 3.3.4] for the characterization of free pro-$p$ groups on sets converging to $1$.
Number Fields
We work with a fixed algebraic closure $\bar{\mathbb{Q}}$ of $\mathbb{Q}$ and a fixed embedding from $\bar{\mathbb{Q}}$ into $\mathbb{C}$. Every number field is considered to be a subfield of $\bar{\mathbb{Q}}$, and hence of $\mathbb{C}$. For an integer $m\geq 1$, we denote the group of $m$-th roots of unity in $\bar{\mathbb{Q}}$ by $\mu_{m}$. Let $F$ be a subfield of $\bar{\mathbb{Q}}$ and $v$ a finite place of $F$. We write
  • $O_{F}$ for the integer ring of $F$,
  • $F_{v}$ for the $v$-adic completion of $F$,
  • $G_{F}$ for the absolute Galois group $\operatorname{Gal}(\bar{\mathbb{Q}}/F)$ of $F$.
Let $K$ be an imaginary quadratic field and $\mathfrak{m}$ a nonzero integral ideal of $K$. We write
  • $K(\mathfrak{m})$ for the ray class field of $K$ of modulo $\mathfrak{m}$. If $\alpha$ generates $\mathfrak{m}$, then we write it as $K(\alpha)$,
  • $K(\mathfrak{m}^{\infty})\coloneqq\cup_{n\geq 1}K(\mathfrak{m}^{n})$.
Elliptic Curves with Complex Multiplication
(cf. Silverman [Sil94, Chapter II])
Let $K$ be an imaginary quadratic field of class number one, and $E$ an elliptic curve over $K$ having complex multiplication by $O_{K}$. For an ideal $\mathfrak{m}$ of $O_{K}$, the $G_{K}$-action on the $\mathfrak{m}$-torsion subgroup $E[\mathfrak{m}]$ determines an injective homomorphism
$\operatorname{Gal}(K(E[\mathfrak{m}])/K)\hookrightarrow\operatorname{Aut}(E[\mathfrak{m}](\bar{\mathbb{Q}}))\cong(O_{K}/\mathfrak{m})^{\times}.$
Moreover, it induces an isomorphism
$\operatorname{Gal}(K(\mathfrak{m})/K)\xrightarrow{\sim}(O_{K}/\mathfrak{m})^{\times}/\operatorname{Im}(O_{K}^{\times}),$